OpenStudy (anonymous):

A juggler is performing an act by juggling several balls. The juggler throws the balls up at an initial height of 3.9 feet with a speed of 14.8 feet per second. If the juggler did not catch a ball, about how long will it take the ball to hit the floor?

OpenStudy (lgbasallote):

im thinking you need to solve for maximum height first. do you know how to do that?

OpenStudy (anonymous):

noo

OpenStudy (lgbasallote):

use the formula v^2 = u^2 + 2ad where v is final speed u is initial speed a is acceleration d is distance <--you're solving for this

OpenStudy (lgbasallote):

so can you tell me the values of v, u and a?

OpenStudy (anonymous):

no i dont understand

OpenStudy (lgbasallote):

which part?

OpenStudy (anonymous):

the whole thing

OpenStudy (lgbasallote):

v^2 = u^2 + 2ad is a formula. do you understand that?